The product of twice a number and three is the same as the difference of five times the n umber an d 3/4. Find the equation and solve.
Answer by RAY100(1637)

Let n= number
.
(2n)(3) = 5n-3/4
.
6n = 5n -3/4
.
n=-3/4,,,,,answer
.
.
check
{2(-3/4)}{3} = 5(-3/4) - 3/4
.
-18/4 = -15/4 -3/4 = -18/4,,,,,,ok
.